# Dual membership certificatesJoonyeob Lee, Dávid Papp, Anita Varga

### v2

@@ 4 Dikin ellipsoids and dual certificates
Throughout this section, we consider again a general proper convex cone $\mathcal{K}$, assuming only that ${({\mathcal{K}^{*}})^{\circ}}$ is the domain of a $\nu$-LHSCB. As before, the gradient and Hessian of this barrier function are denoted by $g$ and $H$, respectively.
@@
Let ${\mathcal{K}^{*}}$ be the domain of a $\nu$-LHSCB, and suppose that $b\in\mathcal{K}$ and $y\in{({\mathcal{K}^{*}})^{\circ}}$ satisfy the relation $b\in\mathscr{B}^{*}(-g(\delta y),r)$ for some $\delta>0$ and $r=\frac{1}{2(\nu+1)}$. Then $y$ is a B-certificate of $b$.
@@
$01+\nu-y^{\mathrm{T}}b\geq 1-\sqrt{\nu}r-\frac{1}{2}r^{2}=1-\frac{\sqrt{\nu}}{2(1+\nu)}-\frac{1}{8(1+\nu)^{2}}\geq\frac{23}{32}>$ 0 .
@@ 5 Computing dual membership certificates
(18) $&\max_{z,\alpha}&&\alpha\\ &\ \textup{s.t.}&&\alpha w+z=b\\ &&&z\in\mathcal{K},\ \alpha\in\mathbb{R}\end{aligned}\qquad\qquad\begin{aligned} &\min_{y}&&b^{\mathrm{T}}y\\ &\ \textup{s.t.}&&w^{\mathrm{T}}y=1\\ &&&y\in\mathcal{K}^{*}\end{aligned}$
It is clear from our assumptions that strong duality holds with attainment on both the primal and dual side, since both problems have a Slater point. The optimal objective function value of (18) is positive if and only if $b\in\mathcal{K}^{\circ}$. However, standard duality theory and an approximately optimal numerical solution of either problem cannot rigorously certify this fact: a strictly feasible dual solution $y$ can be scaled to exactly satisfy the only linear constraint $w^{\mathrm{T}}y=1$, and thus we can obtain a true, verifiable dual feasible solution, however, the corresponding objective value is an upper bound on the exact optimal value, therefore, a nonnegative dual value implies nothing. On the primal side we get a lower bound, and so an exact primal feasible $(z,\alpha)$ with $\alpha\geq 0$ would be a rigorous certificate, but numerical methods for problems of this form (IPAs in particular) tend to yield only approximately feasible primal solutions, meaning that only a vector close to $b$ is certified to be in the cone. In contrast, any $y\in\mathcal{B}(b)$ (or $y\in\mathcal{H}(b)$ if ${\mathcal{K}^{*}}$ is a hyperbolicity cone) does rigorously certify $b\in\mathcal{K}$. Our goal is to compute such a dual certificate $y$.
The problem (18) is a nonsymmetric convex conic optimization problem, with the particular difficulty that no assumptions are made about the cone $\mathcal{K}$. There are only a handful of primal-dual interior-point algorithms (IPAs) developed for this general class of problems, which require only one LHSCB but simultaneously compute near-optimal primal and dual solutions; see, for example, [37, 29, 27, 6]). [^1] In these works it is assumed that the primal cone $\mathcal{K}$ is the domain of a tractable LHSCB, with no assumptions on the dual cone ${\mathcal{K}^{*}}$. We made the opposite assumption for (18) to maintain consistency with the rest of this paper. Since these methods solve the primal and dual problems simultaneously, this is only a notational difference. All of these are suitable tools for numerically addressing the decision problem of whether $b\in\mathcal{K}$. However, if we also aim to obtain dual certificates, it appears that short-step IPAs that operate within a close neighborhood of the central path are the most appropriate methods. This is due to the relationship between the definition of this neighborhood and the sufficient conditions we derived for dual certificates in Theorems 4.1 and 4.3.
@@
(19) $&\inf_{x}&&c^{\mathrm{T}}x\\ &\ \textup{s.t.}&&Ax=b\\ &&&x\in\mathcal{K}\end{aligned}\qquad\qquad\begin{aligned} &\sup_{y}&&b^{\mathrm{T}}y\\ &\ \textup{s.t.}&&A^{\mathrm{T}}y+s=c\\ &&&s\in{\mathcal{K}^{*}}\end{aligned}$
@@
=$\mathcal{F}^{\circ}\overset{\mathrm{def}}{=}\left\{(x,y,s)\in\mathcal{K}^{\circ}\times\mathbb{R}^{m}\times(\mathcal{K}^{*})^{\circ}\,\middle|\,Ax=b,\ A^{\top}y+s=c\right\}.$
@@
where $\eta\in(0,1)$ is the neighborhood radius (a fixed parameter throughout the algorithm, chosen by the user) and $\tau$ is the central path parameter (which tends geometrically to 00 as we approach optimality). Similar definitions appear, for example, in [24, 36, 37].
@@
$\mathcal{F}^{\circ}$ =$=\{(z,\alpha,y)\in\mathcal{K}^{\circ}\times\mathbb{R}\times{({\mathcal{K}^{*}})^{\circ}}\mid\alpha w+z=b,w^{\mathrm{T}}y=1\}\text{ and }$
$\mathcal{N}(\eta,\tau)$ =$=\left\{(z,\alpha,y)\in\mathcal{F}^{\circ}\mid\|y-H(y)^{-1}(z/\tau)\|_{y}\leq\eta\right\}.$
@@ 6 Near-optimal primal feasible solutions
(21) $&\inf_{\gamma\in\mathbb{R}}&&\gamma\\ &\ \textup{s.t.}&&\begin{pmatrix}\gamma\\ -b\end{pmatrix}\in{\mathcal{K}_{c,A}}\end{aligned}\qquad\qquad\begin{aligned} &\sup_{y\in\mathbb{R}^{m}}&&b^{\mathrm{T}}y\\ &\ \textup{s.t.}&&\begin{pmatrix}1\\ y\end{pmatrix}\in{\mathcal{K}^{*}_{c,A}}\end{aligned}$
@@
In addition to our previous assumption that ${({\mathcal{K}^{*}})^{\circ}}$ is the domain of a known $\nu$-LHSCB $f_{\mathcal{K}^{*}}$, let us make the common regularity assumptions that (1) both the primal and dual problems are feasible (therefore we have finite optimal values); (2) $\begin{pmatrix}c^{\mathrm{T}}\\ -A\end{pmatrix}$ has linearly independent rows; and (3) ${\mathcal{K}^{*}_{c,A}}$ has a nonempty interior. The last two assumptions ensure that $({\mathcal{K}^{*}_{c,A}})^{\circ}$ is also the domain of a known $\nu$-LHSCB (Proposition 1.2, Item 4.), namely
@@
Now, suppose that we have an algorithm that we can apply to (21) and compute, using only the dual barrier $f_{{\mathcal{K}^{*}_{c,A}}}$, the (approximately) optimal primal and dual values of the problems (19) and (21), and a (near-)optimal dual feasible solution $y$ for these problems without computing any primal feasible solutions of (19). When $A$ has many more columns than rows, circumventing the iterative optimization of $x$ in this manner can save computational time and memory [28, 27], but many applications call for the actual primal solution to (19), not just the optimal value. As we shall see, we can still apply the same methods, and a primal feasible and (near-)optimal solution to the original problem (19) can be computed in post-processing using our reconstruction formula, at little additional cost. In fact, throughout the algorithms, primal feasible solutions that serve as easily verifiable certificates of our progress towards optimality can be computed in each iteration at a negligible additional cost.
Mirroring our discussion in the Introduction on feasbility certificates, one could argue that these reconstructed feasible solutions are preferable to the numerically computed ones, since they are theoretically guaranteed to be exact if computed (as post-processing) in exact arithmetic. Alternatively, one may use this throughout the algorithm as a guarantee that the numerical method did not veer too far from optimality due to numerical problems (or more reliably detect numerical problems), without resorting to a method that runs entirely in exact or high-precision arithmetic.
@@ 6.1 Exact primal reconstruction in primal-dual IPAs
The arguments for this case are essentially identical to the ones used in Section 5, after noting that (21) can also be equivalently written in the form of (18), with $\gamma$ playing the role of $\alpha$, the vector $e_{0}=(1,0,\dots,0)$ playing the role of $w$, and $(0,-b)$ playing the role of $A$. (The only difference is that $e_{0}$ does not necessarily belong to $({\mathcal{K}_{c,A}})^{\circ}$, but this is irrelevant now.)
With this change in notation, the summary of the previous section is that the application of IPAs using the neighborhood (20) to the optimization problem (21) yields a sequence of iterates $(\gamma,y)$, each of which has the property that $(1,y)$ certifies the relation $(\gamma,-b)\in{\mathcal{K}_{c,A}}$, that is, it certifies the existence of an $x\in\mathcal{K}$ satisfying $Ax=b$ and $c^{\mathrm{T}}x=\gamma$ without ever computing such an $x$ in the interim. We may compute such a primal vector $x$ using our reconstruction theorem (Theorem 3.2), with $f_{{\mathcal{K}^{*}_{c,A}}}$ in place of $f$, $(1,y)$ in place of $y$ and $(\gamma,-b)$ in place of $b$. And since these IPAs require the computation of the gradient and the factorization of the Hessian of the barrier in each iteration, the additional computation represented by the reconstruction formula (13) does not add a substantial computational cost. The bulk of the additional cost is due to matrix-vector multiplications that are cheaper than a single iteration of the IPA.
@@ 6.2 Exact primal reconstruction in dual IPAs
(23) $\min\left\{\tau f_{{\mathcal{K}^{*}_{c,A}}}(y_{0},y)-b^{\mathrm{T}}y\,\middle|\,y_{0}=1,(y_{0},y)\in({\mathcal{K}^{*}_{c,A}})^{\circ}\right\},$
and the neighborhood is defined in terms the local norm associated with the dual barrier. Often, the central path parameter $\tau$ in (23) is equal to or bounded from below by the scaled duality gap $\mu=(c^{\mathrm{T}}x-b^{\mathrm{T}}y)/\nu$ at each iteration, and, crucially, $b^{\mathrm{T}}y_{\tau}$ is within $\nu\tau$ of the optimal objective function value – this is what guarantees linear convergence of the methods in the objective function. See [24, Theorem 5.3.10] for a classic example. We have the following general scheme for primal reconstruction in dual methods that follow the above principle.
@@
Assume that $(1,y)\in({\mathcal{K}^{*}_{c,A}})^{\circ}$ (equivalently, that $(y,s)$ is a dual feasible solution to (19) for some $s\in\mathcal{K}^{\circ}$), and let $\gamma\in\mathbb{R}$ be an arbitrary desired primal objective function value. Denote the gradient and Hessian of the $\nu$-LHSCB $f_{\mathcal{K}^{*}}$ by $g_{{\mathcal{K}^{*}}}$ and $H_{{\mathcal{K}^{*}}}$, and let $B_{{\mathcal{K}^{*}}}=H_{{\mathcal{K}^{*}}}+g_{{\mathcal{K}^{*}}}g_{{\mathcal{K}^{*}}}^{\mathrm{T}}$. Then the vector
(24) $x_{\gamma}\overset{\mathrm{def}}{=}B_{{\mathcal{K}^{*}}}({\upsilon})A_{c}^{\mathrm{T}}\left(A_{c}B_{{\mathcal{K}^{*}}}({\upsilon})A_{c}^{\mathrm{T}}\right)^{-1}b_{\gamma},$
@@
$A_{c}=\begin{pmatrix}c\\ -A\end{pmatrix}\text{, }\quad b_{\gamma}=\begin{pmatrix}\gamma\\ -b\end{pmatrix}\text{, \; and }\;{\upsilon}=A_{c}^{\mathrm{T}}\begin{pmatrix}1\\ y\end{pmatrix}$
@@
As in the rest of this paper, if ${\mathcal{K}^{*}}$ is hyperbolic and $f_{\mathcal{K}^{*}}$ is the corresponding logarithmic barrier, then every instance of $B_{{\mathcal{K}^{*}}}$ in (24) can be replaced by $H_{{\mathcal{K}^{*}}}$.
@@
(25) $-\tau g(1,y_{\tau})=\begin{pmatrix}b^{\mathrm{T}}y_{\tau}+\tau\nu\\ -b\end{pmatrix}.$
Recall from the proof of Theorem 2.6 that Eq. (25) implies that $(1,y_{\tau})$ is both an H- and B-certificate of the membership of the right-hand side vector in ${\mathcal{K}^{*}_{c,A}}$, certifying the existence of a feasible solution $x$ in (19) whose attained value is $c^{\mathrm{T}}x=b^{\mathrm{T}}y_{\tau}+\tau\nu$. From the same theorem, every vector $y$ sufficiently close to $y_{\tau}$ has the same property; that is, $(1,y)$ certifies the existence of a primal feasible solution $x$ whose objective function value is $b^{\mathrm{T}}y+\tau\nu$ [^2] It is possible to make this statement and the remaining “for sufficiently close…” statements of this section more precise and quantitative using Theorem 4.1 and Theorem 4.3, but we omit the details, as some path-following methods may use different neighborhoods anyway.
@@
We can then proceed in one of two ways: for some cones, such as the nonnegative orthant and the positive semidefinite cone, it is possible to find the best certifiable $\gamma$ directly from this affine expression. (For linear programming, this is simple arithmetic, for semidefinite programming it leads to a generalized eigenvalue problem.)
If this is not possible or efficient, we may invoke the sufficient conditions found in Section 4. Comparing Eq. (25) to those results, we see that every vector $(y_{0},y)$ in a suitable Dikin ellipsoid centered at $(1,y_{\tau})$ is a dual certificate of the relation $(b^{\mathrm{T}}y_{\tau}+\tau\nu,-b)\in{\mathcal{K}^{*}_{c,A}}$, certifying that there exists a feasible solution $x$ in (19) whose attained value is $c^{\mathrm{T}}x=b^{\mathrm{T}}y_{\tau}+\tau\nu$. Conversely, $(1,y)$ is a dual certificate of $-g(1,y)$ and also of every vector in the Dikin ellipsoid of radius $r$ around this center. (Here, $r=1$ for H-certificates and $r=(2(\nu+1))^{-1}$ for B-certificates, based on Theorems 4.1 and 4.3.) Thus, for every $y$ sufficiently close to $y_{\tau}$, we have that

### v3

@@ 4 Dikin ellipsoids and dual certificates
Throughout this section, we consider again a general proper convex cone $\mathcal{K}$, assuming only that $({\mathcal{K}^{*}})^{\circ}$} is the domain of a $\nu$-LHSCB.
@@
Let $({\mathcal{K}^{*}})^{\circ}$} be the domain of a $\nu$-LHSCB, and suppose that $b\in\mathcal{K}$ and $y\in{({\mathcal{K}^{*}})^{\circ}}$ satisfy the relation $b\in\mathscr{B}^{*}(-g(\delta y),r)$ for some $\delta>0$ and $r=\frac{1}{2(\nu+1)}$. Then $y$ is a B-certificate of $b$.
@@
$1+\nu-y^{\mathrm{T}}b\geq 1-\sqrt{\nu}r-\frac{1}{2}r^{2}=1-\frac{\sqrt{\nu}}{2(1+\nu)}-\frac{1}{8(1+\nu)^{2}}\geq\frac{23}{32}>0.$
@@ 5 Computing dual membership certificates
(18) $\begin{aligned} &\max_{z,\alpha}&&\alpha\\ &\ \textup{s.t.}&&\alpha w+z=b\\ &&&z\in\mathcal{K},\ \alpha\in\mathbb{R}\end{aligned}\qquad\qquad\begin{aligned} &\min_{y}&&b^{\mathrm{T}}y\\ &\ \textup{s.t.}&&w^{\mathrm{T}}y=1\\ &&&y\in\mathcal{K}^{*}\end{aligned}$
It is clear from our assumptions that strong duality holds with attainment on both the primal and dual side, since both problems have a Slater point. The optimal objective function value $\alpha^{*}$ of (18) is positive if and only if $b\in\mathcal{K}^{\circ}$. However, standard duality theory and an approximately optimal numerical solution of either problem cannot rigorously certify this fact: a strictly feasible dual solution $y$ can be scaled to exactly satisfy the only linear constraint $w^{\mathrm{T}}y=1$, and thus we can obtain a true, verifiable dual feasible solution, however, the corresponding objective value is an upper bound on the exact optimal value, therefore, a nonnegative dual value implies nothing. On the primal side we get a lower bound, and so an exact primal feasible $(z,\alpha)$ with $\alpha\geq 0$ would be a rigorous certificate, but numerical methods for problems of this form (interior-point algorithms in particular) tend to yield only approximately feasible primal solutions, meaning that only a vector close to $b$ is certified to be in the cone. In contrast, any $y\in\mathcal{B}(b)$ (or $y\in\mathcal{H}(b)$ if $\mathcal{K}^{*}$} is a hyperbolicity cone) does rigorously certify $b\in\mathcal{K}$. Our goal is to compute such a dual certificate $y$.
The problem (18) is a nonsymmetric convex conic optimization problem, with the particular difficulty that no assumptions are made about the cone $\mathcal{K}$. There are only a handful of primal-dual interior-point algorithms (IPAs) developed for this general class of problems, which require only one LHSCB but simultaneously compute near-optimal primal and dual solutions; see, for example, [39, 31, 29, 6]). [^1] In these works it is assumed that the primal cone $\mathcal{K}$ admits a tractable LHSCB, with no assumptions on the dual cone $\mathcal{K}^{*}$}. We made the opposite assumption for (18) to maintain consistency with the rest of this paper. Since these methods solve the primal and dual problems simultaneously, this is only a notational difference. All of these are suitable tools for numerically addressing the decision problem of whether $b\in\mathcal{K}$. However, if we also aim to obtain dual certificates, it appears that short-step IPAs that operate within a close neighborhood of the central path are the most appropriate methods. This is due to the relationship between the definition of this neighborhood and the sufficient conditions we derived for dual certificates in Theorems 4.1 and 4.3.
@@
(19) $\begin{aligned} &\inf_{x}&&c^{\mathrm{T}}x\\ &\ \textup{s.t.}&&Ax=b\\ &&&x\in\mathcal{K}\end{aligned}\qquad\qquad\begin{aligned} &\sup_{y}&&b^{\mathrm{T}}y\\ &\ \textup{s.t.}&&A^{\mathrm{T}}y+s=c\\ &&&s\in{\mathcal{K}^{*}}\end{aligned}$
@@
$\mathcal{F}^{\circ}\overset{\mathrm{def}}{=}\left\{(x,y,s)\in\mathcal{K}^{\circ}\times\mathbb{R}^{m}\times(\mathcal{K}^{*})^{\circ}\,\middle|\,Ax=b,\ A^{\mathrm{T}}y+s=c\right\}.$
@@
where $\eta\in(0,1)$ is the neighborhood radius (a fixed parameter throughout the algorithm, chosen by the user) and $\tau$ is the central path parameter (which tends geometrically to 0 as we approach optimality). Similar neighborhoods are used, for example, in [26, 38, 39].
@@
$\mathcal{F}^{\circ}$ $=\{(z,\alpha,y)\in\mathcal{K}^{\circ}\times\mathbb{R}\times{({\mathcal{K}^{*}})^{\circ}}\mid\alpha w+z=b,w^{\mathrm{T}}y=1\}\text{ and }$
$\mathcal{N}(\eta,\tau)$ $=\left\{(z,\alpha,y)\in\mathcal{F}^{\circ}\mid\|y-H(y)^{-1}(z/\tau)\|_{y}\leq\eta\right\}.$
@@ 6 Near-optimal primal feasible solutions
(21) $\begin{aligned} &\inf_{\gamma\in\mathbb{R}}&&\gamma\\ &\ \textup{s.t.}&&\begin{pmatrix}\gamma\\ -b\end{pmatrix}\in{\mathcal{K}_{c,A}}\end{aligned}\qquad\qquad\begin{aligned} &\sup_{y\in\mathbb{R}^{m}}&&b^{\mathrm{T}}y\\ &\ \textup{s.t.}&&\begin{pmatrix}1\\ y\end{pmatrix}\in{\mathcal{K}^{*}_{c,A}}\end{aligned}$
@@
In addition to our previous assumption that $({\mathcal{K}^{*}})^{\circ}$} is the domain of a known $\nu$-LHSCB $f_{\mathcal{K}^{*}}$, let us make following common regularity assumptions:
  1. 1)
    Both the primal and dual problems are feasible (therefore we have finite optimal values);
  2. 2)
    $\begin{pmatrix}c^{\mathrm{T}}\\ -A\end{pmatrix}$ has linearly independent rows; and
  3. 3)
    $\mathcal{K}^{*}_{c,A}$} has a nonempty interior.
The last two assumptions ensure that $({\mathcal{K}^{*}_{c,A}})^{\circ}$ is also the domain of a known $\nu$-LHSCB (Proposition 1.2, Item 4.), namely
@@
Now, suppose that we have an algorithm that we can apply to (21) and compute, using only the dual barrier $f_{{\mathcal{K}^{*}_{c,A}}}$, the (approximately) optimal primal and dual values and a (near-)optimal dual feasible solution $y$ for this problem. As the optimal objective values and the dual feasible regions of (21) and (19) are the same, this also solves the dual problem of (19) without computing any of its primal feasible solutions. When $A$ has many more columns than rows, circumventing the iterative optimization of $x$ in this manner can save computational time and memory [30, 29]. On the other hand, many applications call for the actual primal solution to (19), not just the optimal value. As we shall see, in that case, we can still apply the same methods, and a primal feasible and (near-)optimal solution to the original problem (19) can be computed in post-processing using our reconstruction formula, at little additional cost. In fact, throughout the algorithms, primal feasible solutions that serve as easily verifiable certificates of our progress towards optimality can be computed in each iteration at a negligible additional cost.
Mirroring our discussion in the Introduction on feasbility certificates, one could argue that these reconstructed feasible solutions are preferable to the numerically computed ones, since they are theoretically guaranteed to be exact if computed (as post-processing) in higher-precision or exact arithmetic. Alternatively, one may use this throughout the algorithm as a guarantee that the numerical method did not veer too far from optimality due to numerical problems (or more reliably detect numerical problems), without resorting to a method that runs entirely in exact or high-precision arithmetic.
@@ 6.1 Exact primal reconstruction in primal-dual IPAs
The arguments for this case are essentially identical to the ones used in Section 5, after noting that (21) can also be equivalently written in the form of (18), with $\gamma$ playing the role of $\alpha$, the vector $e_{0}=(1,0,\dots,0)$ playing the role of $w$, and (0,b)(0,-b) playing the role of $A$. (The only difference is that $e_{0}$ does not necessarily belong to $({\mathcal{K}_{c,A}})^{\circ}$, but this is irrelevant now.)
With this change in notation, the summary of the previous section is that the application of IPAs using the neighborhood (20) to the optimization problem (21) yields a sequence of iterates $(\gamma,y)$, each of which has the property that $(1,y)$ certifies the relation $(\gamma,-b)\in{\mathcal{K}_{c,A}}$, that is, it certifies the existence of an $x\in\mathcal{K}$ satisfying $Ax=b$ and $c^{\mathrm{T}}x=\gamma$ without ever computing such an $x$ in the process. We may compute such a primal vector $x$ using our reconstruction theorem (Theorem 3.2), with $f_{{\mathcal{K}^{*}_{c,A}}}$ in place of $f$, $(1,y)$ in place of $y$ and $(\gamma,-b)$ in place of $b$. And since these IPAs require the computation of the gradient and the factorization of the Hessian of the barrier in each iteration, the additional computation represented by the reconstruction formula (13) does not add a substantial computational cost. The bulk of the additional cost is due to matrix-vector multiplications that are cheaper than a single iteration of the IPA.
@@ 6.2 Exact primal reconstruction in dual IPAs
(23) $y_{\tau}\overset{\mathrm{def}}{=}\arg\min\left\{\tau f_{{\mathcal{K}^{*}_{c,A}}}(y_{0},y)-b^{\mathrm{T}}y\,\middle|\,y_{0}=1,(y_{0},y)\in({\mathcal{K}^{*}_{c,A}})^{\circ}\right\},$
and the neighborhood is defined in terms the local norm associated with the dual barrier. Often, the central path parameter $\tau$ in (23) is equal to or bounded from below by the scaled duality gap $\mu=(c^{\mathrm{T}}x-b^{\mathrm{T}}y)/\nu$ at each iteration, and, crucially, $b^{\mathrm{T}}y_{\tau}$ is within $\nu\tau$ of the optimal objective function value – this is what guarantees linear convergence of the methods in the objective function. (See [26, Theorem 5.3.10] for a classic example.) We have the following general scheme for primal reconstruction in dual methods that follow the above principle.
@@
Assume that $(1,y)\in({\mathcal{K}^{*}_{c,A}})^{\circ}$ (equivalently, that $(y,s)$ is a dual feasible solution to (19) for some $s\in\mathcal{K}^{\circ}$), and let $\gamma\in\mathbb{R}$ be an arbitrary scalar. Denote the Hessian of the $\nu$-LHSCB $f_{\mathcal{K}^{*}}$ by $H_{{\mathcal{K}^{*}}}$. Then the vector
(24) $x_{\gamma}\overset{\mathrm{def}}{=}H_{{\mathcal{K}^{*}}}({\upsilon})A_{c}^{\mathrm{T}}\left(A_{c}H_{{\mathcal{K}^{*}}}({\upsilon})A_{c}^{\mathrm{T}}\right)^{-1}b_{\gamma},$
@@
$A_{c}=\begin{pmatrix}c^{\mathrm{T}}\\ -A\end{pmatrix}\text{, }\quad b_{\gamma}=\begin{pmatrix}\gamma\\ -b\end{pmatrix}\text{, \; and }\;{\upsilon}=A_{c}^{\mathrm{T}}\begin{pmatrix}1\\ y\end{pmatrix}$
@@
(25) $\tau g(1,y_{\tau})=\begin{pmatrix}b^{\mathrm{T}}y_{\tau}+\tau\nu\\ -b\end{pmatrix}$.
Recall from the proof of Theorem 2.6 that Eq. (25) implies that $(1,y_{\tau})$ is both an H- and B-certificate of the membership of the right-hand side vector in $\mathcal{K}^{*}_{c,A}$}, certifying the existence of a feasible solution $x$ in (19) whose attained value is $c^{\mathrm{T}}x=b^{\mathrm{T}}y_{\tau}+\tau\nu$. From the same theorem, every vector $y$ sufficiently close to $y_{\tau}$ has the same property; that is, $(1,y)$ certifies the existence of a primal feasible solution $x$ whose objective function value is $b^{\mathrm{T}}y+\tau\nu$. [^2] It is possible to make this statement and the remaining “for sufficiently close…” statements of this section more precise and quantitative using Theorem 4.1 and Theorem 4.3, but we omit the details, as some path-following methods may use different neighborhoods anyway.
@@
We can then proceed in one of two ways: for some cones, such as the nonnegative orthant and the positive semidefinite cone, it is possible to find the best certifiable $\gamma$ directly from this affine expression. (For linear programming this is simple arithmetic, for semidefinite programming it leads to a generalized eigenvalue problem.)
If this is not possible or efficient enough, we may invoke the sufficient conditions found in Section 4. Comparing Eq. (25) to those results, we see that every vector $(y_{0},y)$ in a suitable Dikin ellipsoid centered at $(1,y_{\tau})$ is a dual certificate of the relation $(b^{\mathrm{T}}y_{\tau}+\tau\nu,-b)\in{\mathcal{K}^{*}_{c,A}}$, certifying that there exists a feasible solution $x$ in (19) whose attained value is $c^{\mathrm{T}}x=b^{\mathrm{T}}y_{\tau}+\tau\nu$. Conversely, $(1,y)$ is a dual certificate of g(1,y)-g(1,y) and also of every vector in the Dikin ellipsoid of radius $r$ around this center. (Here, $r=1$ for H-certificates and $r=(2(\nu+1))^{-1}$ for B-certificates, based on Theorems 4.1 and 4.3.) Thus, for every $y$ sufficiently close to $y_{\tau}$, we have that